# MQED-QD: An Open-Source Package for Quantum Dynamics Simulation in Complex Dielectric Environments

### 05 Mar 2026 v1

@@ 1 Introduction
Exciton dynamics of molecular assemblies at the interface of plasmonic nanostructures has emerged as a central theme in organic optoelectronics, 43, 7, 47, 51, 75, 50, 1 light harvesting, 59, 9, 29, 74, 45, 53, 58 and quantum information science.57, 42, 30, 69, 32, 17 In such systems, the interaction between molecular excitation and the electromagnetic environment can profoundly modify energy transfer dynamics. For example, nearby metallic nanoparticles and thin dielectric films, resonance energy transfer (RET) can be enhanced or suppressed by orders of magnitude beyond the vacuum Förster limit 23, 2, 35, 14 because the photonic density of states near a dielectric or metallic interface reshapes the dipole–dipole coupling strength that mediates exciton transport. 20, 71, 49, 27, 13 Furthermore, molecular polaritonics experiments have further shown that strong light–matter coupling in optical cavities creates hybrid exciton–photon states with dramatically altered transport properties.72, 73, 25, 31, 67, 33 From theory and simulation perspectives, the key challenge is to treat the dynamics of a large ensemble of quantum emitters and their interactions with the dielectric environment on an equal footing.
@@
The macroscopic quantum electrodynamics (MQED) theory, developed by Welsch and co-workers, offers an elegant framework by quantizing the electromagnetic field in the presence of arbitrary, dispersive, and absorbing media through the fluctuation–dissipation theorem.56, 3 In the MQED framework, all radiative inter-molecular coupling, dissipation rates, and energy level shifts can be expressed in terms of the classical dyadic Green’s function, which encodes the full influence of the dielectric environment. This feature has been applied to generalize Förster’s $1/R^{6}$ energy transfer theory to include retardation, material dispersion, and geometric effects of the surrounding medium,20 and elucidate polariton-coupled electron transfer under the influence of dielectric environment. 12 Despite much progress, these results almost always rely on high-symmetry geometries, in which the dyadic Green’s function can be obtained analytically, such as Sommerfeld integrals for planar interfaces or Mie theory for spheres. 49, 39, 70, 44 For a nanostructure of arbitrary geometry or dispersion relation, one must resort to numerical methods such as the boundary element method (BEM) 34, the finite element method (FEM)40 , and the finite-difference time-domain (FDTD) method.65
@@
Beyond the calculation of parameters related to electromagnetic fields, a second major challenge lies in integrating the MQED framework with the quantum dynamics of molecular emitters. The conventional approach utilizes semiclassical simulations that self-consistently combine Maxwell’s equations with Ehrenfest molecular dynamics38, 8, 46 to capture rotation-vibration effects at complex plasmonic interfaces. 63, 62, 64 While recent quantum-corrected models have extended these classical simulations to incorporate nonlocal and tunneling effects,21, 55 the full MQED treatment of the field often leads to complex quantum dynamics problems characterized by long-range coupling, non-Markovian memory kernels, and generalized dissipation.11 Capturing these dynamical effects accurately requires sophisticated methods, such as generalized quantum master equations or hierarchical equations of motion.4, 15Consequently, there remains an outstanding need for a platform that bridges numerically accurate Green’s functions with efficient dynamical solvers for the open quantum system description of molecular excitons.
@@
In this work, we developed MQED-QD, an open-source Python package to (i) extract the dyadic Green’s function from classical electromagnetic solvers, (ii) construct the MQED Hamiltonian and the corresponding quantum master equation, and (iii) propagate the quantum dynamics for coupled molecular emitters. The MQED-QD package allows for user-specified geometry and material parameters and facilitates the full pipeline for calculating the dyadic Green’s functions by analytical solutions or numerical methods with a built-in calibration protocol. Through interfacing with QuTiP,41 the MQED-QD package simulates the dynamics of open quantum systems by integrating either the standard Lindblad master equation or a non-Hermitian Schrödinger equation. Another key feature of the MQED-QD package is that the entire workflow is accessible through a set of command-line tools, lowering the barrier for non-specialist users. For demonstration, we investigate exciton transport and delocalization properties of molecular aggregates in the presence of silver nanostructures of various geometries, including a planar surface and a nanorod.
@@ 2 Theoretical Framework
In this section, we provide an overview of the macroscopic quantum electrodynamics (MQED) and the quantum master equation approach for the dynamics of an ensemble of quantum emitters. For more detailed derivation and discussion, the interested reader is referred to Ref. 36.
@@ 2.1 The MQED Hamiltonian
Here $\mathbf{\hat{f}^{\dagger}}(\mathbf{r},\omega)$ and $\mathbf{\hat{f}}(\mathbf{r},\omega)$ are the creation and annihilation operators of the quantized bosonic vector fields derived from the quantization of the hybrid medium excitations and photons,28, 18 rather than photons in a vacuum. By design, the vector components of these quantized bosonic field operator satisfy the bosonic commutation relation, ${\hat{f}_{i}}(\mathbf{r},\omega),{\hat{f}_{j}^{\dagger}}(\mathbf{r}^{\prime},\omega^{\prime})]=\delta_{ij}\delta(\mathbf{r}-\mathbf{r}^{\prime})\delta(\omega-\omega^{\prime})$ for $i,j\in\{x,y,z\}$. Namely, these vector field operators correspond to create and annihilate photons dressed by dielectric medium.
In the multipolar gauge,22 the coupling between the quantum subsystems and the dressed photon field can be expressed in the form
@@
The transition dipole moment of the $\alpha$-th subsystem is $\boldsymbol{\mu}_{\alpha}=\mu_{\alpha}\mathbf{n}_{\alpha}$ where $\mu_{\alpha}$ is its real-valued magnitude and $\mathbf{n}_{\alpha}$ is the unit vector of its direction. The electric field operator at position $\mathbf{r}_{\alpha}$ takes the form20:
@@ 2.2 Quantum Dynamics of Molecular Exciton States
To investigate the dynamics of the quantum subsystems interacting with the quantized electromagnetic fields, we then treat the hybrid system within the framework of open quantum systems. Starting from the Heisenberg equation of motion, we can partially trace out the continuum photonic degrees of freedom (photonic bath) and focus on the dynamics of reduced density matrix $\hat{\rho}_{\mathrm{M}}(t)=\mathrm{Tr}_{\mathrm{P}}[\hat{\rho}(t)]=\int d^{3}\mathbf{r}\int_{0}^{\infty}d\omega\hat{\rho}(\mathbf{r},\omega,t)$ where we assume the initial condition of the photonic bath is in the vacuum state. The Heisenberg equation can be reduced to a master equation with the generalized spectral density 54, 11, 10:
@@
In this work, we focus on the weak light-matter coupling regime and assume that the generalized spectral density varies smoothly near the molecular transition frequency $\omega_{\text{M}}$. Thus, we can evoke the Markov approximation and eliminate the spectral dependence of $J_{\alpha\beta}(\omega)$ by taking the two-point dyadic Green’s function as $\overline{\overline{\mathbf{G}}}(\mathbf{r}_{\alpha},\mathbf{r}_{\beta},\omega_{\mathrm{M}})$. Namely, the electromagnetic environment is considered a broadband Markovian reservoir with negligible memory effects. Under this Markovian approximation, the resulting quantum master equation for the reduced density matrix takes the Lindbladian form10, 66
@@
where $\hat{O}_{1},\hat{O}_{2}]=\hat{O}_{1}\hat{O}_{2}-\hat{O}_{2}\hat{O}_{1}$ and $\{\hat{O}_{1},\hat{O}_{2}\}=\hat{O}_{1}\hat{O}_{2}+\hat{O}_{2}\hat{O}_{1}$ are the commutator and the anti-commutator of two operators, respectively. Here, $\hat{H}_{\mathrm{CP}}$ and $\hat{H}_{\mathrm{DDI}}$ represent the Casimir-Polder (CP) potential and the dipole-dipole interactions (DDI), respectively11, 10, 66.
@@
and the CP potential $\Lambda_{\alpha}^{\mathrm{Sc}}$ describes the transition energy shift induced by the dielectric environment, which can be evaluated by:11, 66
@@
Note that, in the weak coupling limit, the diagonal term of the dissipation matrix $\Gamma_{\alpha\alpha}$ describes the population decay of the electronic excited state of a single molecule in a complex dielectric environment.19. The ratio between the dissipation rate ($\Gamma_{\alpha\alpha}$) and the spontaneous emission rate ($\Gamma_{0}$) corresponds to the Purcell enhancement factor $\Gamma_{\alpha\alpha}/\Gamma_{0}=\frac{6\pi c}{\omega_{\mathrm{M}}}\mathbf{n}_{\alpha}\cdot\mathrm{Im}\overline{\overline{\mathbf{G}}}(\mathbf{r}_{\alpha},\mathbf{r}_{\alpha},\omega_{\mathrm{M}})\cdot\mathbf{n}_{\alpha}$.52 Thus, the single-point dyadic Green’s function can be evaluated in terms of the Purcell factor, which is available in classical electrodynamics simulations.
@@ 2.3 Exciton Transport and Delocalization
Within this MQED framework, we characterize exciton transport and delocalization using two key observables: the mean-square displacement (MSD) and the participation ratio (PR).48, 60 Given a single excitation initially localized at site $\alpha_{0}$, the MSD quantifies the spatial spread of the exciton from the initial site, while the PR measures the effective number of sites participating in the quantum state. The formal definitions of MSD and PR, including the renormalization procedure required for dissipative dynamics, are presented in Section 4. In the following sections, we demonstrate the robustness of the MQED-QD package by benchmarking the MSD and PR in complex dielectric environments. Note that, since we integrate the quantum master equation for the time evolution of the reduced density matrix, the current framework is not limited to these observables.
@@ 3.1 Analytical Solutions
For specific geometric structures, such as an isotropic sphere and an infinite planar surface, we can employ the analytical solutions to Maxwell’s equations in terms of spectral expansion. The scattering of electromagnetic waves by an isotropic, spherical particle can be evaluated as a sum of spherical harmonic functions using Mie theory,39 which can be used for calculating the dyadic Green’s functions.44 For a single point source over a planar layered dielectrics, one can expand the spherical wave in terms of sum of plane waves and evanescent waves using Fourier transformation in space, known as the Sommerfeld identity.37 The reflection component of the dyadic Green’s function can be directly evaluated using the Fresnel reflection coefficients.70
@@ 3.2 Numerical Methods
To construct the dyadic Green’s function, we utilize the fundamental relationship between an oscillating source point dipole and the electric fields it generates61, 36, 16.
@@ 3.2.1 Boundary Element Method (BEM)
We employ the boundary element method (BEM)24, specifically the MNPBEM toolbox designed for the simulation of metallic nanoparticles (MNP).34 In contrast to general volume-based simulation tools, such as FDTD, the BEM approach discretizes material interfaces and reformulates Maxwell’s equations into surface integral equations. The dielectric environment is then described through local and isotropic dielectric functions $\epsilon_{\lambda}(\omega)$ within each domain (labeled by $\lambda$) separated by well-defined boundaries. Consequently, by restricting discretization to the interfaces rather than the entire volume, the BEM significantly reduces the degrees of freedom and the computational cost compared to volume-based methods.
@@ 4.1 Integrating Lindblad Master Equation
We can reduce the generalized Lindblad master equation (eq 9) to the standard Lindblad form for the density matrix propagation. First, to avoid numerical issues, we enforce the dissipation matrix to be positive semidefinite (i.e. ensuring all eigenvalues are non-negative and real values) by symmetrization:68 $\bar{\Gamma}_{\alpha\beta}=\frac{1}{2}({\Gamma}_{\alpha\beta}+{\Gamma}_{\beta\alpha}^{*})$. Next, we diagonalize the symmetrized matrix $\bar{\Gamma}_{\alpha\beta}]$ to obtain eigenvalues $\gamma_{\alpha}$ and the unitary transformation $U_{\alpha\beta}$ (i.e. $\bar{\Gamma}_{\alpha\beta}=\sum_{\alpha^{\prime}}U_{\alpha\alpha^{\prime}}\gamma_{\alpha^{\prime}}U_{\beta\alpha^{\prime}}^{*}$), so that we can define the jump operators by $\hat{L}_{\alpha}=\sum_{\alpha^{\prime}}U_{\alpha^{\prime}\alpha}^{*}\hat{\sigma}_{\alpha^{\prime}}=\sum_{\alpha^{\prime}}U_{\alpha^{\prime}\alpha}^{*}|\mathrm{G}\rangle\langle\mathrm{X}_{\alpha^{\prime}}|$. In the end, the standard form of the Lindblad master equation reads
@@ 4.2 Non-Hermitian Hamiltonian Approach
Since we focus on the dynamics of excitation states, we can employ an effective non-Hermitian Hamiltonian approach to further simplify the calculation.5, 26, 6 Instead of propagating the reduced density matrix in eq 9, we consider the state vector of the quantum subsystems restricted in the single-excitation manifold, $|\Psi(t)\rangle=\sum_{\alpha=1}^{N_{\mathrm{M}}}c_{\alpha}(t)|\mathrm{X}_{\alpha}\rangle$, following the Schrödinger equation $i\hbar\frac{\partial}{\partial t}|\Psi(t)\rangle=\hat{H}_{\mathrm{eff}}|\Psi(t)\rangle$ with the effective non-Hermitian Hamiltonian defined as

### 06 Mar 2026 v2

@@ 1 Introduction
Exciton dynamics of molecular assemblies at the interface of plasmonic nanostructures has emerged as a central theme in organic optoelectronics, 44, 8, 48, 52, 76, 51, 2 light harvesting, 60, 10, 30, 75, 46, 54, 59 and quantum information science.58, 43, 31, 70, 33, 18 In such systems, the interaction between molecular excitation and the electromagnetic environment can profoundly modify energy transfer dynamics. For example, nearby metallic nanoparticles and thin dielectric films, resonance energy transfer (RET) can be enhanced or suppressed by orders of magnitude beyond the vacuum Förster limit 24, 3, 36, 15 because the photonic density of states near a dielectric or metallic interface reshapes the dipole–dipole coupling strength that mediates exciton transport. 21, 72, 50, 28, 14 Furthermore, molecular polaritonics experiments have further shown that strong light–matter coupling in optical cavities creates hybrid exciton–photon states with dramatically altered transport properties.73, 74, 26, 32, 68, 34 From theory and simulation perspectives, the key challenge is to treat the dynamics of a large ensemble of quantum emitters and their interactions with the dielectric environment on an equal footing.
@@
The macroscopic quantum electrodynamics (MQED) theory, developed by Welsch and co-workers, offers an elegant framework by quantizing the electromagnetic field in the presence of arbitrary, dispersive, and absorbing media through the fluctuation–dissipation theorem.57, 4 In the MQED framework, all radiative inter-molecular coupling, dissipation rates, and energy level shifts can be expressed in terms of the classical dyadic Green’s function, which encodes the full influence of the dielectric environment. This feature has been applied to generalize Förster’s $1/R^{6}$ energy transfer theory to include retardation, material dispersion, and geometric effects of the surrounding medium,21 and elucidate polariton-coupled electron transfer under the influence of dielectric environment. 13 Despite much progress, these results almost always rely on high-symmetry geometries, in which the dyadic Green’s function can be obtained analytically, such as Sommerfeld integrals for planar interfaces or Mie theory for spheres. 50, 40, 71, 45 For a nanostructure of arbitrary geometry or dispersion relation, one must resort to numerical methods such as the boundary element method (BEM) 35, the finite element method (FEM)41 , and the finite-difference time-domain (FDTD) method.66
@@
Beyond the calculation of parameters related to electromagnetic fields, a second major challenge lies in integrating the MQED framework with the quantum dynamics of molecular emitters. The conventional approach utilizes semiclassical simulations that self-consistently combine Maxwell’s equations with Ehrenfest molecular dynamics39, 9, 47 to capture rotation-vibration effects at complex plasmonic interfaces. 64, 63, 65 While recent quantum-corrected models have extended these classical simulations to incorporate nonlocal and tunneling effects,22, 56 the full MQED treatment of the field often leads to complex quantum dynamics problems characterized by long-range coupling, non-Markovian memory kernels, and generalized dissipation.12 Capturing these dynamical effects accurately requires sophisticated methods, such as generalized quantum master equations or hierarchical equations of motion.5, 16Consequently, there remains an outstanding need for a platform that bridges numerically accurate Green’s functions with efficient dynamical solvers for the open quantum system description of molecular excitons.
@@
In this work, we developed MQED-QD, an open-source Python package to (i) extract the dyadic Green’s function from classical electromagnetic solvers, (ii) construct the MQED Hamiltonian and the corresponding quantum master equation, and (iii) propagate the quantum dynamics for coupled molecular emitters. The MQED-QD package allows for user-specified geometry and material parameters and facilitates the full pipeline for calculating the dyadic Green’s functions by analytical solutions or numerical methods with a built-in calibration protocol. Through interfacing with QuTiP,42 the MQED-QD package simulates the dynamics of open quantum systems by integrating either the standard Lindblad master equation or a non-Hermitian Schrödinger equation. Another key feature of the MQED-QD package is that the entire workflow is accessible through a set of command-line tools, lowering the barrier for non-specialist users. For demonstration, we investigate exciton transport and delocalization properties of molecular aggregates in the presence of silver nanostructures of various geometries, including a planar surface and a nanorod.
@@ 2 Theoretical Framework
In this section, we provide an overview of the macroscopic quantum electrodynamics (MQED) and the quantum master equation approach for the dynamics of an ensemble of quantum emitters. For more detailed derivation and discussion, the interested reader is referred to Ref. 37.
@@ 2.1 The MQED Hamiltonian
Here $\mathbf{\hat{f}^{\dagger}}(\mathbf{r},\omega)$ and $\mathbf{\hat{f}}(\mathbf{r},\omega)$ are the creation and annihilation operators of the quantized bosonic vector fields derived from the quantization of the hybrid medium excitations and photons,29, 19 rather than photons in a vacuum. By design, the vector components of these quantized bosonic field operator satisfy the bosonic commutation relation, ${\hat{f}_{i}}(\mathbf{r},\omega),{\hat{f}_{j}^{\dagger}}(\mathbf{r}^{\prime},\omega^{\prime})]=\delta_{ij}\delta(\mathbf{r}-\mathbf{r}^{\prime})\delta(\omega-\omega^{\prime})$ for $i,j\in\{x,y,z\}$. Namely, these vector field operators correspond to create and annihilate photons dressed by dielectric medium.
In the multipolar gauge,23 the coupling between the quantum subsystems and the dressed photon field can be expressed in the form
@@
The transition dipole moment of the $\alpha$-th subsystem is $\boldsymbol{\mu}_{\alpha}=\mu_{\alpha}\mathbf{n}_{\alpha}$ where $\mu_{\alpha}$ is its real-valued magnitude and $\mathbf{n}_{\alpha}$ is the unit vector of its direction. The electric field operator at position $\mathbf{r}_{\alpha}$ takes the form21:
@@ 2.2 Quantum Dynamics of Molecular Exciton States
To investigate the dynamics of the quantum subsystems interacting with the quantized electromagnetic fields, we then treat the hybrid system within the framework of open quantum systems. Starting from the Heisenberg equation of motion, we can partially trace out the continuum photonic degrees of freedom (photonic bath) and focus on the dynamics of reduced density matrix $\hat{\rho}_{\mathrm{M}}(t)=\mathrm{Tr}_{\mathrm{P}}[\hat{\rho}(t)]=\int d^{3}\mathbf{r}\int_{0}^{\infty}d\omega\hat{\rho}(\mathbf{r},\omega,t)$ where we assume the initial condition of the photonic bath is in the vacuum state. The Heisenberg equation can be reduced to a master equation with the generalized spectral density 55, 12, 11:
@@
In this work, we focus on the weak light-matter coupling regime and assume that the generalized spectral density varies smoothly near the molecular transition frequency $\omega_{\text{M}}$. Thus, we can evoke the Markov approximation and eliminate the spectral dependence of $J_{\alpha\beta}(\omega)$ by taking the two-point dyadic Green’s function as $\overline{\overline{\mathbf{G}}}(\mathbf{r}_{\alpha},\mathbf{r}_{\beta},\omega_{\mathrm{M}})$. Namely, the electromagnetic environment is considered a broadband Markovian reservoir with negligible memory effects. Under this Markovian approximation, the resulting quantum master equation for the reduced density matrix takes the Lindbladian form11, 67
@@
where $\hat{O}_{1},\hat{O}_{2}]=\hat{O}_{1}\hat{O}_{2}-\hat{O}_{2}\hat{O}_{1}$ and $\{\hat{O}_{1},\hat{O}_{2}\}=\hat{O}_{1}\hat{O}_{2}+\hat{O}_{2}\hat{O}_{1}$ are the commutator and the anti-commutator of two operators, respectively. Here, $\hat{H}_{\mathrm{CP}}$ and $\hat{H}_{\mathrm{DDI}}$ represent the Casimir-Polder (CP) potential and the dipole-dipole interactions (DDI), respectively12, 11, 67.
@@
and the CP potential $\Lambda_{\alpha}^{\mathrm{Sc}}$ describes the transition energy shift induced by the dielectric environment, which can be evaluated by:12, 67
@@
Note that, in the weak coupling limit, the diagonal term of the dissipation matrix $\Gamma_{\alpha\alpha}$ describes the population decay of the electronic excited state of a single molecule in a complex dielectric environment.20. The ratio between the dissipation rate ($\Gamma_{\alpha\alpha}$) and the spontaneous emission rate ($\Gamma_{0}$) corresponds to the Purcell enhancement factor $\Gamma_{\alpha\alpha}/\Gamma_{0}=\frac{6\pi c}{\omega_{\mathrm{M}}}\mathbf{n}_{\alpha}\cdot\mathrm{Im}\overline{\overline{\mathbf{G}}}(\mathbf{r}_{\alpha},\mathbf{r}_{\alpha},\omega_{\mathrm{M}})\cdot\mathbf{n}_{\alpha}$.53 Thus, the single-point dyadic Green’s function can be evaluated in terms of the Purcell factor, which is available in classical electrodynamics simulations.
@@ 2.3 Exciton Transport and Delocalization
Within this MQED framework, we characterize exciton transport and delocalization using two key observables: the mean-square displacement (MSD) and the participation ratio (PR).49, 61 Given a single excitation initially localized at site $\alpha_{0}$, the MSD quantifies the spatial spread of the exciton from the initial site, while the PR measures the effective number of sites participating in the quantum state. The formal definitions of MSD and PR, including the renormalization procedure required for dissipative dynamics, are presented in Section 4. In the following sections, we demonstrate the robustness of the MQED-QD package by benchmarking the MSD and PR in complex dielectric environments. Note that, since we integrate the quantum master equation for the time evolution of the reduced density matrix, the current framework is not limited to these observables.
@@ 3.1 Analytical Solutions
For specific geometric structures, such as an isotropic sphere and an infinite planar surface, we can employ the analytical solutions to Maxwell’s equations in terms of spectral expansion. The scattering of electromagnetic waves by an isotropic, spherical particle can be evaluated as a sum of spherical harmonic functions using Mie theory,40 which can be used for calculating the dyadic Green’s functions.45 For a single point source over a planar layered dielectrics, one can expand the spherical wave in terms of sum of plane waves and evanescent waves using Fourier transformation in space, known as the Sommerfeld identity.38 The reflection component of the dyadic Green’s function can be directly evaluated using the Fresnel reflection coefficients.71
@@ 3.2 Numerical Methods
To construct the dyadic Green’s function, we utilize the fundamental relationship between an oscillating source point dipole and the electric fields it generates62, 37, 17.
@@ 3.2.1 Boundary Element Method (BEM)
We employ the boundary element method (BEM)25, specifically the MNPBEM toolbox designed for the simulation of metallic nanoparticles (MNP).35 In contrast to general volume-based simulation tools, such as FDTD, the BEM approach discretizes material interfaces and reformulates Maxwell’s equations into surface integral equations. The dielectric environment is then described through local and isotropic dielectric functions $\epsilon_{\lambda}(\omega)$ within each domain (labeled by $\lambda$) separated by well-defined boundaries. Consequently, by restricting discretization to the interfaces rather than the entire volume, the BEM significantly reduces the degrees of freedom and the computational cost compared to volume-based methods.
@@ 4.1 Integrating Lindblad Master Equation
We can reduce the generalized Lindblad master equation (eq 9) to the standard Lindblad form for the density matrix propagation. First, to avoid numerical issues, we enforce the dissipation matrix to be positive semidefinite (i.e. ensuring all eigenvalues are non-negative and real values) by symmetrization:69 $\bar{\Gamma}_{\alpha\beta}=\frac{1}{2}({\Gamma}_{\alpha\beta}+{\Gamma}_{\beta\alpha}^{*})$. Next, we diagonalize the symmetrized matrix $\bar{\Gamma}_{\alpha\beta}]$ to obtain eigenvalues $\gamma_{\alpha}$ and the unitary transformation $U_{\alpha\beta}$ (i.e. $\bar{\Gamma}_{\alpha\beta}=\sum_{\alpha^{\prime}}U_{\alpha\alpha^{\prime}}\gamma_{\alpha^{\prime}}U_{\beta\alpha^{\prime}}^{*}$), so that we can define the jump operators by $\hat{L}_{\alpha}=\sum_{\alpha^{\prime}}U_{\alpha^{\prime}\alpha}^{*}\hat{\sigma}_{\alpha^{\prime}}=\sum_{\alpha^{\prime}}U_{\alpha^{\prime}\alpha}^{*}|\mathrm{G}\rangle\langle\mathrm{X}_{\alpha^{\prime}}|$. In the end, the standard form of the Lindblad master equation reads
@@ 4.2 Non-Hermitian Hamiltonian Approach
Since we focus on the dynamics of excitation states, we can employ an effective non-Hermitian Hamiltonian approach to further simplify the calculation.6, 27, 7 Instead of propagating the reduced density matrix in eq 9, we consider the state vector of the quantum subsystems restricted in the single-excitation manifold, $|\Psi(t)\rangle=\sum_{\alpha=1}^{N_{\mathrm{M}}}c_{\alpha}(t)|\mathrm{X}_{\alpha}\rangle$, following the Schrödinger equation $i\hbar\frac{\partial}{\partial t}|\Psi(t)\rangle=\hat{H}_{\mathrm{eff}}|\Psi(t)\rangle$ with the effective non-Hermitian Hamiltonian defined as